OpenStudy (anonymous):

iF(X+Y)=0.17 and coxX=0.50, find sinY .

OpenStudy (anonymous):

You can find x from the second equation. Now you've got x. Stick that into the first equation to get y. Then get the sin of y from 3rd.

OpenStudy (anonymous):

i get x = 60

OpenStudy (anonymous):

so if i put it into the cos(60+y)=0.17 hows that?

OpenStudy (anonymous):

Are you sure that that is all there is to the problem? 2 things seem to be wrong with the question. 1) is cox supposed to be cos? and 2) is your trig in radians or degrees?

OpenStudy (anonymous):

1) that is cos. typo error 2) it doesn't have a degree sign. it's just a number. but maybe it's a radian

OpenStudy (anonymous):

Whoever worded that problem did a really poor job. You can TRY to assume it's radians but that's just an assumption and it isn't necessarily logical. I can't help you with the problem stated that way.

OpenStudy (anonymous):

anyway, when you substitute 60 to cos(x+y) = 0.17 how can i solve?

OpenStudy (anonymous):

Is it cos(x+y)=0.17 or (x+y)=0.17 like in your original question? Now the question changed.

OpenStudy (anonymous):

i'm sorry. typo error again. it's cos(x+y)= 0.17.

OpenStudy (anonymous):

how can i solve it? :)

OpenStudy (anonymous):

ok 2 things.

OpenStudy (anonymous):

you know cosX = 0.5, so you can get sinX. Now, use cos(a + b) = cosa x cosb - sina x sinb. You know that this equals a numerical amount since that is given and you have cosX, so you can isolate sinY and solve. This is all you have to do. Jusy work the formulas.

OpenStudy (anonymous):

i got 0.3455 is it correct? :)

OpenStudy (anonymous):

You got it!! Good job!

OpenStudy (anonymous):

thanks a lot friend ;)

OpenStudy (anonymous):

you're welcome.